# Interpreting the Weight Space of Customized Diffusion Models

### v2

@@ Abstract
We investigate the space of weights spanned by a large collection of customized diffusion models. We populate this space by creating a dataset of over 60,000 models, each of which is a base model fine-tuned to insert a different person’s visual identity. We model the underlying manifold of these weights as a subspace, which we term weights2weights. We demonstrate three immediate applications of this space – sampling, editing, and inversion. First, as each point in the space corresponds to an identity, sampling a set of weights from it results in a model encoding a novel identity. Next, we find linear directions in this space corresponding to semantic edits of the identity (e.g., adding a beard). These edits persist in appearance across generated samples. Finally, we show that inverting a single image into this space reconstructs a realistic identity, even if the input image is out of distribution (e.g., a painting). Our results indicate that the weight space of fine-tuned diffusion models behaves as an interpretable latent space of identities. [^1] Project page: https://snap-research.github.io/weights2weights
@@
Figure 1: weights2weights (w2w) space enables generative applications on model weights. We model a manifold of customized diffusion model weights as a subspace encoding different human identities. This forms a space that supports inverting the visual identity from a single image into a model, editing the identity encoded in the model, and sampling new models that encode diverse instances of people. This space effectively serves as a latent space over customized models.
@@ 1 Introduction
Generative models have emerged as a powerful tool to model our rich visual world. In particular, the latent space of single-step generative models, such as Generative Adversarial Networks (GANs) [15, 26], has been shown to linearly encode meaningful concepts in the output images. For instance, linear directions in GANs encode different attributes (e.g., gender or age of faces) and can be composed for multi-attribute image edits [56, 17, 22]. Alas, in multi-step generative models, like diffusion models [57, 19], such a linear latent space is yet to be found.
@@
Recently introduced personalization approaches, such as Dreambooth [50] or Custom Diffusion [31], may hint at where such an interpretable latent space can exist in diffusion models. These methods aim to learn an instance of a subject, such as a person’s visual identity. Rather than searching for a latent code that represents an identity in the input noise space, these approaches customize diffusion models by fine-tuning on subject images, which results in identity-specific model weights. We therefore hypothesize that a latent space can exist in the weights themselves.
@@
To test our hypothesis, we fine-tune over 60,000 personalized models on individual identities to obtain points that lie on a manifold of customized diffusion model weights. To reduce the dimensionality of each data point, we use low-rank approximation (LoRA) [20] during fine-tuning and further apply Principal Components Analysis (PCA) to the set of data points. This forms our final space: weights2weights (w2w). Unlike GANs, which model the pixel space of images, we model the weight space of these personalized models. Thus, each sample in our space corresponds to an identity-specific model which can consistently generate that subject. We provide a schematic in Fig. 2 that contrasts the GAN latent space with our proposed w2w space, demonstrating the differences and analogies between these two representations.
@@
Creating this space unlocks a variety of applications that involve traversal in w2w (Fig. 1). First, we demonstrate that sampling model weights from w2w space corresponds to a new identity. Second, we find linear directions in this space corresponding to semantic edits of the identity. Finally, we show that enforcing weights to live in this space enables a diffusion model to learn an identity given a single image, even if it is out of distribution.
@@
We find that w2w space is highly expressive through quantitative evaluation on editing customized models and encoding new identities given a single image. Qualitatively, we observe this space supports sampling models that encode diverse and realistic identities, while also capturing the key characteristics of out-of-distribution identities.
@@
Figure 2: The weights2weights space enables GAN-like applications on identity-encoding model weights. Novel identities can be sampled from the space and edited by linearly traversing along semantic directions in weight space. Additionally, a single image can be inverted into the space to produce a model that consistently generates that identity.
@@ 2 Related Work
Latent space of generative models. Linear latent space models of facial shape and appearance were studied extensively in the 1990s, using both PCA-based representations (e.g. Active Appearance Models [7], 3D Morphable Models [4]) as well as models operating directly in pixel and keypoint space [49]. However, these techniques were restricted to aligned and cropped frontal faces. More recently, Generative Adversarial Networks (GANs), particularly the StyleGAN series [25, 26, 27, 24], have showcased editing capabilities facilitated by their interpretable latent space. Furthermore, linear directions can be found in their latent space to conduct semantic edits by training linear classifiers or applying PCA [56, 17], among other methods for discovering semantic directions [60, 6]. Several methods aim to project real images into the GAN latent space in order to conduct this editing [69, 1, 47, 59].
Although diffusion models architecturally lack such a latent space, some works aim to discover a GAN-like latent space in them. This has been explored in the UNet bottleneck layer [32, 37], noise space [8], and text-embedding space [3]. Concept Sliders [14] explores the weight space for semantic image editing by conducting low-rank training with contrasting image or text pairs.
@@ 3 Method
We start by demonstrating how we create a manifold of model weights as illustrated in Fig. 3. We explain how we obtain low-dimensional data points for this space, each of which represents an individual identity. We then use these points to model a weights manifold. Next, we find linear directions in this manifold that correspond to semantic attributes and use them for editing the identities. Finally, we demonstrate how this manifold can be utilized for constraining an ill-posed inversion task with a single image to reconstruct its identity.
@@ 3.1 Preliminaries
where the first term corresponds to the standard diffusion denoising objective using the subject-specific data $\mathbf{x}$ conditioned on the text prompt “[identifier] [class noun]” (e.g., “[v] person”), denoted $\mathbf{c}$. The second term, weighted by $\lambda$, corresponds to a prior preservation loss, which involves the standard denoising objective using the model’s own generated samples $\mathbf{x}^{\prime}$ for the broader class $\mathbf{c}^{\prime}$ (e.g., “person”). This prevents the model from associating the class name with the specific instance, while also leveraging the semantic prior on the class. We utilize this approach to obtain a per-subject model and use its weights to create the interpretable weights manifold.
@@ 3.2 Constructing the weights manifold
Modeling the weights manifold. We posit that our data $D\subseteq\mathbb{R}^{d}$ lies on a lower-dimensional manifold of weights that encode identities. A randomly sampled set of weights in $\mathbb{R}^{d}$, would not be guaranteed to produce a valid model encoding identity as the $d$ degrees of freedom can be fine-tuned for any purpose. Therefore, we hypothesize that this manifold is a subset of the weight space. Inspired by findings that high-level concepts can be encoded as linear subspaces of representations [34, 44, 41], we model this subset as a linear subspace $\mathbb{R}^{m}$ where $m<d$, and call it weights2weights (w2w) space. We represent points in this subspace as a linear combination of basis vectors $\textbf{w}=\{w_{1},...,w_{m}\}$, $w_{i}\in\mathbb{R}^{d}$. In practice, we apply Principal Component Analysis (PCA) on the $N$ models and keep the first $m$ principal components for dimensional reduction and forming our basis of $m$ vectors.
@@ 3.3 Finding Interpretable Weight Space Directions
We seek a direction $\mathbf{n}\in\mathbb{R}^{d}$ defining a hyperplane that separates between binary identity properties embedded in the model weights (e.g., male/female), similarly to hyperplanes observed in the latent space of GANs [56]. We assume binary labels are given for attributes present in the identities encoded by the models. We then train linear classifiers using weights of the models as data based on these labels, imposing separating hyperplanes in weight space. Given an identity parameterized by weights $\theta$, we can manipulate a single attribute by traversing in a direction $\mathbf{n}$, orthogonal to the separating hyperplane: $\theta_{\text{edit}}=\theta+\alpha{\mathbf{n}}$.
@@ 3.4 Inversion into w2w Space
Traditionally, inversion of a generative model involves finding an input such as a latent code that best reconstructs a given image [64, 35]. This corresponds to finding a projection of the input onto the learned data manifold [69]. With w2w space, we model a manifold of data which happens to be model weights rather than images. Inspired by latent optimization methods [1, 69], we propose a gradient-based method of inverting a single identity from an image into our discovered space.
@@ 4 Experiments
We demonstrate w2w space on human identities for a variety of applications. We begin by detailing implementation details. Next, we use w2w space for 1) sampling new models encoding novel identities, 2) editing identity attributes in a consistent manner via linear traversal in w2w space, 3) embedding a new identity given a single image, and 4) projecting out-of-distribution identities into w2w space. Finally, we analyze how scaling the number of models in our dataset of model weights affects the disentanglement of attribute directions and preservation of identity. We refer the reader to Appendix H, where we apply w2w space to other visual concepts.
@@ 4.2 Sampling from w2w Space
We present images generated from models that were sampled from the weights manifold (i.e., w2w Space) in Fig. 4. We follow the sampling procedure from Sec. 3.2, and generate images from the sampled model with various prompts and seeds. As shown, each new model encodes a novel, realistic, and consistent identity. Additionally, we present the nearest neighbor model among the training dataset of model weights. We use cosine similarity on the models’ principal component representations. Comparing with the nearest neighbors shows that these samples are not just copies from the dataset, but rather encode diverse identities with different attributes. Yet, the samples still demonstrate some similar features to the nearest neighbors. These include jawline and eye shape (top row), facial hair (middle row), and nose and eye shape (bottom row). Appendix A includes more such examples.
@@ 4.3 Editing Subjects
w2w edits are identity preserving and disentangled. We evaluate over a range of identity-specific attributes and present three (gender, chubby, narrow eyes) in Tab. 1. Edits in w2w preserve the identity of the original subject as measured by the ID score. These edits are semantically aligned with the desired effect as indicated by the CLIP score while minimally interfering with other visual concepts, as measured by LPIPS. We note that the CLIP score can be noisy in this setting as text captions can be too coarse to describe attributes as nuanced as those related to the human face.
@@
Composing edits. Edit directions in w2w space can be composed linearly as shown in Fig. 6. The composed edits persist in appearance across different generations, binding to the identity. Furthermore, the edited weights result in a new model, where the subject has different attributes while still maintaining as much of the prior identity. As we operate on a weight manifold, minimal changes are made to other concepts, such as scene layout or other people. For instance, in Fig. 6, adding edits to the woman does not interfere with Obama standing by her.
@@
Figure 6: Composing multiple attributes persist in appearance. We demonstrate how we can apply multiple edits in w2w space without significant degradation of the original identity or interference with other concepts. These edits persist and maintain their appearance across different generation seeds and prompts.
@@ 4.4 Inverting Subjects
We compare our results to two approaches that use Dreambooth with rank-1 LoRA. The first is trained on a single image. The second is trained on multiple images of each identity. We generate such images by following our identity dataset construction from Sec. 4.1). This approach can be viewed as a pseudo-upper bound on modeling identity as it uses multiple images.
w2w space provides a strong identity prior. Inverting a single image into w2w space improves on the single image Dreambooth baseline and closes the gap with the Dreambooth baseline that uses multiple identity images (Tab. 2). Conducting Dreambooth fine-tuning with a single image in the original weight space leads to image overfitting and poor subject reconstruction as indicated by a lower ID score. In contrast, by constraining the optimized weights to lie on a manifold of identity weights, w2w inversion inherits the rich priors of the models used to discover the space. As such, it can extract a high-fidelity identity that is consistent and compositional across generations. We present qualitative comparisons against Dreambooth and single-image Dreambooth in Appendix C.
@@
Inverted models are editable. Fig. 7 demonstrates that a diverse set of identities can be faithfully represented in w2w space. After inversion, the encoded identity can be composed in novel contexts and poses. For instance, the inverted man (rightmost example) can be seen posing with Taylor Swift or rendered as a statue. Moreover, semantic edits can be applied to the inverted models while maintaining appearance across generations.
Figure 7: Single image inversion reconstructs identity and enables editing in w2w space. Inverted identities can be composed in novel contexts. Additionally, applying our discovered semantic directions edits attributes that persist in appearance across generation seeds and prompts.
@@ 4.6 Effect of Number of Models Spanning w2w Space
Figure 8: Projecting out-of-distribution identities. We show that our inversion method can convert unrealistic identities into realistic renderings with in-domain facial features. The resulting identity can be composed in novel scenes, such as playing tennis or rendered into other artistic domains.
Figure 9: Scaling our model weights further disentangles classifier directions. We highlight the trend in disentanglement of three examples where attributes may be strongly correlated among identities. As the number of models (Identities) is increased, the features are less entangled.

### v3

@@ Abstract
We investigate the space of weights spanned by a large collection of customized diffusion models. We populate this space by creating a dataset of over 60,000 models, each of which is a base model fine-tuned to insert a different person’s visual identity. We model the underlying manifold of these weights as a subspace, which we term weights2weights. We demonstrate three immediate applications of this space that result in new diffusion models – sampling, editing, and inversion. First, sampling a set of weights from this space results in a new model encoding a novel identity. Next, we find linear directions in this space corresponding to semantic edits of the identity (e.g., adding a beard), resulting in a new model with the original identity edited. Finally, we show that inverting a single image into this space encodes a realistic identity into a model, even if the input image is out of distribution (e.g., a painting). We further find that these linear properties of the diffusion model weight space extend to other visual concepts. Our results indicate that the weight space of fine-tuned diffusion models can behave as an interpretable meta-latent space producing new models. [^1] Project page: https://snap-research.github.io/weights2weights
@@
Figure 1: weights2weights (w2w) space enables controllable creation of new customized diffusion models. We model a manifold of customized diffusion models as a subspace of weights that encodes different instances of a broad visual concept (e.g., human identities, dog breeds, etc.). This forms a space that supports inverting the subject (e.g., identity) from a single image into a model, editing the subject encoded in the model, and sampling new models that encode new instances of the visual concept. Each of these operations results in a new model that can consistently generate the subject.
@@ 1 Introduction
Generative models have emerged as a powerful tool to model our rich visual world. In particular, the latent space of single-step generative models, such as generative adversarial networks (GANs) [18, 28], has been shown to linearly encode meaningful concepts in the output images. For instance, datasets of latent vectors were used to discover linear directions in the GAN latent space encoding different attributes (e.g., gender or age of faces) [60, 20]. Even earlier, datasets of images and keypoints were leveraged to discover subspaces of facial shape and appearance [6, 53].
@@
We aim to extend this even further, using datasets of model weights instead datasets of images or latents. Can we discover such interpretable subspaces in the model weights themselves? Recently introduced personalization approaches, such as Dreambooth [54] or Custom Diffusion [34], may hint that this is the case. These methods aim to learn an instance of a subject, such as a person’s visual identity. Rather than searching for a latent code that represents an identity in the input noise space, these approaches customize diffusion models by fine-tuning on subject-specific images, which results in identity-specific model weights. We therefore hypothesize that a latent space can exist in the weights themselves.
@@
To test our hypothesis, we fine-tune over 60,000 personalized models on individual identities to obtain points that lie on a manifold of customized diffusion model weights. To reduce the dimensionality of each data point, we use low-rank approximation (LoRA) [23] during fine-tuning and further apply Principal Components Analysis (PCA) to the set of data points. This forms our final space: weights2weights (w2w). Unlike traditional generative models like GANs, which model the pixel space of images, we model the weight space of these personalized models. Thus, each sample in our space corresponds to an identity-specific model which can consistently generate that subject. We provide a schematic in Fig. 2 that contrasts a typical latent space with our proposed w2w space, demonstrating the differences and analogies between these two representations. w2w space can be thought of as a meta-latent space, enabling controllable creation of new models instead of just images like a traditional latent space.
@@
Creating this space unlocks a variety of applications that involve traversal in w2w (Fig. 1). First, we demonstrate that sampling model weights from w2w space corresponds to a new model encoding a novel subject. Second, we find linear directions in this space corresponding to semantic edits of the identity. Finally, we show that enforcing weights to live in this space enables a diffusion model to learn a subject given a single image, even if it is out of distribution.
@@
We find that w2w space is highly expressive through quantitative evaluation on editing customized models and encoding new identities given a single image. Qualitatively, we observe this space supports sampling models that encode diverse and realistic identities, while also capturing the key characteristics of out-of-distribution identities. We finally demonstrate that similar weight subspaces exist for other visual concepts such as dog breeds and car types.
@@
Figure 2: The weights2weights space operates as a meta-latent space. Unlike a traditional generative latent space, w2w space controls the model itself rather than single image instances. New identity-encoding models can be sampled from the space and edited by linearly traversing along semantic directions in weight space. Additionally, a single image can be inverted into the space to produce a model that consistently generates that identity.
@@ 2 Related Work
Latent space of generative models. Linear latent space models of facial shape and appearance were studied extensively in the 1990s, using PCA-based representations (e.g. Active Appearance Models [9], 3D Morphable Models [6]) as well as operating directly in pixel and keypoint space [53]. However, these techniques were restricted to aligned and cropped frontal faces. More recently, generative adversarial networks (GANs), particularly the StyleGAN series [27, 28, 29, 26], have showcased editing capabilities facilitated by their interpretable latent space. Furthermore, linear directions can be found in their latent space to conduct semantic edits by training linear classifiers or applying PCA [60, 20], among other methods for discovering semantic directions [66, 8]. Several methods aim to project real images into the GAN latent space in order to conduct this editing [77, 3, 1, 51, 64]. Beyond the latent space, works such as [4] found that directions could be discovered in the neuron activation space, suggesting the interpretability of weights.
Although diffusion models architecturally lack a GAN-like latent space, some works aim to discover similar spaces in these models. This has been explored in the UNet bottleneck layer [35, 41], noise space [78, 10], and text-embedding space [5]. Concept Sliders [17] explores the weight space for semantic image editing by conducting low-rank training with contrasting image or text pairs.
@@ 3 Method
We start by demonstrating how we create a manifold of model weights as illustrated in Fig. 3. We explain how we obtain low-dimensional data points for this space, each of which represents an individual subject from a broad class (i.e., identity). We then use these points to model a weights manifold. Next, we find linear directions in this manifold that correspond to semantic attributes and use them for editing the identities. Finally, we demonstrate how this manifold can be utilized for constraining an ill-posed inversion task with a single image to reconstruct its identity.
@@ 3.1 Preliminaries
where the first term corresponds to the standard diffusion denoising objective using the subject-specific data $\mathbf{x}$ conditioned on the text prompt “[identifier] [class noun]” (e.g., “[v] person”), denoted $\mathbf{c}$. The second term, weighted by $\lambda$, corresponds to a prior preservation loss, which involves the standard denoising objective using the model’s own generated samples $\mathbf{x}^{\prime}$ for the broader class $\mathbf{c}^{\prime}$ (e.g., “person”). This prevents the model from associating the class name with the specific instance, while also leveraging the semantic prior on the class.
@@ 3.2 Constructing the weights manifold
Modeling the weights manifold. We posit that our data $D\subseteq\mathbb{R}^{d}$ lies on a lower-dimensional manifold of weights that encode identities. A randomly sampled set of weights in $\mathbb{R}^{d}$, would not be guaranteed to produce a valid model encoding identity as the $d$ degrees of freedom can be fine-tuned for any purpose. Therefore, we hypothesize that this manifold is a subset of the weight space. Inspired by findings that high-level concepts can be encoded as linear subspaces of representations [37, 48, 45, 13], we model this subset as a linear subspace $\mathbb{R}^{m}$ where $m<d$, and call it weights2weights (w2w) space. We represent points in this subspace as a linear combination of basis vectors $\textbf{w}=\{w_{1},...,w_{m}\}$, $w_{i}\in\mathbb{R}^{d}$. In practice, we apply Principal Component Analysis (PCA) on the $N$ models and keep the first $m$ principal components for dimensional reduction and forming our basis of $m$ vectors.
@@ 3.3 Finding Interpretable Weight Space Directions
We seek a direction $\mathbf{n}\in\mathbb{R}^{d}$ defining a hyperplane that separates between binary identity properties embedded in the model weights (e.g., male/female), similarly to hyperplanes observed in the latent space of GANs [60]. We assume binary labels are given for attributes present in the identities encoded by the models. We then train linear classifiers using weights of the models as data based on these labels, imposing separating hyperplanes in weight space. Given an identity parameterized by weights $\theta$, we can manipulate a single attribute by traversing in a direction $\mathbf{n}$, orthogonal to the separating hyperplane: $\theta_{\text{edit}}=\theta+\alpha{\mathbf{n}}$. An edit operation in w2w spaee produces a new model with the original subject edited, allowing the model to generate infinitely many new images of the edited subject.
@@ 3.4 Inversion into w2w Space
Traditionally, inversion of a generative model involves finding an input such as a latent code that best reconstructs a given image [70, 38]. This corresponds to finding a projection of the input onto the learned data manifold [77]. With w2w space, we model a manifold of model weights rather than images. Inspired by latent optimization methods [1, 77], we propose a gradient-based method of inverting a single identity from an image into our discovered space.
@@ 4 Experiments
We demonstrate w2w space on the visual concept of human identities for a variety of applications. We begin by detailing implementation details. Next, we use w2w space for 1) sampling new models encoding novel identities, 2) editing identity attributes in a consistent manner via linear traversal in w2w space, 3) embedding a new identity given a single image, and 4) projecting out-of-distribution identities into w2w space. Finally, we analyze how scaling the number of models in our dataset of model weights affects the disentanglement of attribute directions and preservation of identity.
@@ 4.2 Sampling from w2w Space
We present images generated from models that were sampled from the weights manifold (i.e., w2w Space) in Fig. 4. We follow the sampling procedure from Sec. 3.2, and generate images from the sampled model. As shown, each new model encodes a novel, realistic, and consistent identity. Additionally, we present the nearest neighbor model among the training dataset of model weights. We use cosine similarity on the models’ principal component representations. Comparing with the nearest neighbors shows that these samples are not just copies from the dataset, but rather encode diverse identities with different attributes. Yet, the samples still demonstrate some similar features to the nearest neighbors. These include jawline and eye shape (top row), facial hair (middle row), and nose and eye shape (bottom row). Appendix A includes more such examples.
@@ 4.3 Editing Subjects
w2w edits are identity preserving and disentangled. We evaluate over a range of identity-specific attributes and present three (gender, chubby, narrow eyes) in Tab. 1. Edits in w2w preserve the identity of the original subject as measured by the ID score. These edits are semantically aligned with the desired effect as indicated by the CLIP score while minimally interfering with other visual concepts, as measured by LPIPS. We note that the CLIP score can be noisy in this setting as text captions can be too coarse to describe attributes as nuanced as those related to the human face. We supplement this with a user study presented in Appendix B.
@@
Composing edits. Edit directions in w2w space can be composed linearly as shown in Fig. 6. The first column represents samples from the original model, and each subsequent column represents samples from the edited models. Each row shares the same fixed random generation seed. The composed edits persist in appearance across different generations, binding to the identity. Furthermore, the edited weights result in a new model, where the subject has different attributes while still maintaining as much of the prior identity. This is in contrast to editing in a traditional latent space, where an edit only corresponds to a single image. Additionally, as we operate on an personalized identity-specific weight manifold, minimal changes are made to other concepts, such as scene layout or other people. For instance, in Fig. 6, adding edits to the woman does not interfere with the person standing by her.
@@
Figure 6: Composing edits in w2w space. Each column represents fixed seed samples from an edited model. Multiple edits in w2w space minimally degrade the original identity or interfere with other concepts, while maintaining edit appearance across different samples.
@@ 4.4 Inverting Subjects
We compare our results to two approaches that use Dreambooth with rank-1 LoRA. The first is trained on a single image. The second is trained on multiple images of each identity. We generate such images by following our identity dataset construction from Sec. 4.1. This approach can be viewed as a pseudo-upper bound on modeling identity as it uses multiple images.
w2w space provides a strong identity prior. Inverting a single image into w2w space improves on the single image Dreambooth baseline and closes the gap with the Dreambooth baseline that uses multiple identity images (Tab. 2). Conducting Dreambooth fine-tuning with a single image in the original weight space leads to image overfitting and poor subject reconstruction as indicated by a lower ID score. In contrast, by constraining the optimized weights to lie on a manifold of identity weights, w2w inversion inherits the rich priors of the models used to discover the space. As such, it can extract a high-fidelity identity that is consistent and compositional across generations. We present qualitative comparisons against Dreambooth and single-image Dreambooth in Appendix C. We additionally compare against other personalization methods in that section.
@@
Inverted models are editable. Fig. 7 demonstrates that a diverse set of identities can be faithfully represented in w2w space. After inversion, the encoded identity can be composed in novel contexts and poses. For instance, the inverted man (rightmost example) can be seen posing with a celebrity or rendered as a statue. Moreover, semantic edits can be applied to the inverted models while maintaining appearance across generations.
Figure 7: Single image inversion reconstructs identity and enables editing in w2w space. We present generated samples from the inverted models. These inverted identities can be composed in novel contexts and edited using our discovered semantic directions in weight space. These edits persist in appearance across generation seeds and prompts.
Figure 8: Projecting out-of-distribution identities. We show that our inversion method can convert unrealistic identities into realistic renderings with in-domain facial features. Each image represents a generated sample from the inverted model. The resulting identities can be composed in novel scenes, such as playing tennis or rendered into other artistic domains.
@@ 4.6 Effect of Number of Models Spanning w2w Space
Figure 9: Scaling dataset of models further disentangles classifier directions. We highlight the trend in disentanglement of three examples where attributes may be strongly correlated among identities. As the number of models is increased, the features are less entangled.
@@
5 Extending to Other Domains
We extend our hypothesis of interpretable linear weight subspaces in diffusion models to other visual concepts beyond human identities. We apply the weights2weights framework to form a subspace for models encoding different dog breeds. To create a dataset for fine-tuning, we generate images with Stable Diffusion based on each of the $120$ dog classes from ImageNet [11]. We then conduct Dreambooth fine-tuning on each set of dog breed images to create a dataset of $120$ dog-encoding models, subsequently applying PCA. To find edit directions, we use GPT-4 [2] to create labels for each dog breed (e.g., wavy hair or not) and then train linear classifiers on the model weight principal component projections like Sec. 3.3. We apply the same framework to different car categories using models fine-tuned on images from a dataset of 197 different car types [33]. We present results for traversing edit directions in these two subspaces in Fig. 11. Each column represents samples from an edited model. Each row shares the same fixed random generation seed.
Our results provide further evidence that diffusion models can encode visual concepts linearly. This enables the creation of new models in a controlled manner via simple interpolation. For instance, in Fig. 11, we rewrite the model’s learned concept of a small golden retriever to make it bigger, or the model’s encoding of a red van to make it a sports car. Additionally, unlike older PCA-based methods [53, 6, 61, 65] which rely on aligned pixels or keypoints of human faces, weights2weights can extend to other domains beyond human identities. We refer the reader to Appendix G for more results of applying w2w space to other visual concepts.
Figure 11: weights2weights linear subspaces can be created for other visual concepts. We follow the same procedure of applying PCA and finding edit directions with linear classifiers on datasets of models encoding dog breeds and models encoding car types.